OpenStudy (anonymous):

You spin a spinner, and the probability that the spinner will land on red is 18%. Suppose you spin the spinner 4 times. What is the probability that the spinner will land on red at least 2 times? Round to the nearest tenth, if necessary

OpenStudy (anonymous):

10.1% 11.9% 24.3% 63.8%

OpenStudy (perl):

ok lets do this

OpenStudy (kropot72):

The can be solved by using the binomial distribution: \[P(2\ red)=\left(\begin{matrix}4 \\ 2\end{matrix}\right)0.18^{2}\times 0.82^{2}=0.1307\] \[P(3\ red)=\left(\begin{matrix}4 \\ 3\end{matrix}\right)0.18^{3}\times 0.82^{1}=0.0191\] \[P(4\ red)=\left(\begin{matrix}4 \\ 4\end{matrix}\right)0.18^{4}\times 0.82^{0}=0.00105\] Adding the above three values of probability will give the required value.

OpenStudy (anonymous):

add which one ?

OpenStudy (kropot72):

Required probability is 0.1307 + 0.0191 + 0.00105 = ?

OpenStudy (anonymous):

@kropot72 0.15

OpenStudy (kropot72):

BTW none of the answer choices that you posted is correct.

OpenStudy (anonymous):

that's what it says im not sure

OpenStudy (kropot72):

The probability as a percentage is 15.1% rounded to the nearest tenth.